### v1

# SIV-Bench: A Video Benchmark for Social Interaction Understanding and Reasoning

Abstract
The rich and multifaceted nature of human social interaction, encompassing multimodal cues, unobservable relations and mental states, and dynamical behavior, presents a formidable challenge for artificial intelligence. To advance research in this area, we introduce SIV-Bench, a novel video benchmark for rigorously evaluating the capabilities of Multimodal Large Language Models (MLLMs) across Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP). SIV-Bench features 2,792 video clips and 8,792 meticulously generated question-answer pairs derived from a human-LLM collaborative pipeline. It is originally collected from TikTok and YouTube, covering a wide range of video genres, presentation styles, and linguistic and cultural backgrounds. It also includes a dedicated setup for analyzing the impact of different textual cues—original on-screen text, added dialogue, or no text. Our comprehensive experiments on leading MLLMs reveal that while models adeptly handle SSU, they significantly struggle with SSR and SDP, where Relation Inference (RI) is an acute bottleneck, as further examined in our analysis. Our study also confirms the critical role of transcribed dialogue in aiding comprehension of complex social interactions. By systematically identifying current MLLMs’ strengths and limitations, SIV-Bench offers crucial insights to steer the development of more socially intelligent AI. The dataset and code are available at https://kfq20.github.io/sivbench/.
1 Introduction
The development of Multimodal Large Language Models (MLLMs), which can process text, images, and video, marks a significant step toward more human-like AI [46, 53, 64, 20, 55]. These models have shown strong performance across a range of tasks, driving progress in areas such as visual reasoning, video captioning, and multimodal dialogue. As these capabilities expand, there is a growing need for benchmarks that can evaluate model performance, uncover limitations, and guide future research [13, 11, 61, 37, 19]. One critical yet underexplored area is the understanding and reasoning about social interactions — a core aspect of social intelligence that encompasses not only observable behaviors but also implicit mental states and social relationships governing behaviors such as forming bonds, exchanging information, and coordinating actions [4, 42]. However, current video benchmarks—whether designed for specific tasks like object segmentation [17, 54], captioning [36], and activity recognition [60], or for broader video understanding across various domains [24, 25]—still lack a detailed focus on social interactions.
We define the capacity of MLLMs to understand and reason social interactions in video through three core, interrelated dimensions: 1) Social Scene Understanding (SSU) is foundational, enabling the recognition of visible elements such as objects, environments, and socially salient human features like body movements, clothing, and physical appearance. Without reliable scene perception, models cannot anchor their interpretations in contextually relevant cues. 2) Social State Reasoning (SSR) is essential for interpreting the unobservable states of interaction, such as emotions, intents, attitudes, and interpersonal relationships, which guide and shape behavior [44, 52]. This capacity allows models to move beyond surface-level features and grasp the underlying states. 3) Social Dynamics Prediction (SDP) enables the model to reason about how interactions proceed over time or under alternative conditions, capabilities essential for a flexible and human-like understanding of social scenarios [39, 6]. It involves factual prediction and counterfactual prediction. The former anticipates upcoming actions or emotional shifts. The latter examines how changes in social relations or behaviors might alter the outcome of an interaction.
In this work, we introduce SIV-Bench (Social Interaction Video Benchmark), a novel benchmark designed to evaluate MLLMs across the above three core dimensions. SIV-Bench is fundamentally organized around social relationships, recognizing their critical role in shaping social interaction [47, 5, 16]. Specifically, SIV-Bench is built on Fiske’s Relational Models Theory [12], categorizing social interactions via four foundational models (Communal Sharing, Authority Ranking, Equality Matching, and Market Pricing), which are instantiated through 14 specific relation types (e.g., parent-child, friends, colleagues). This relational schema underpins fine-grained evaluation across three key dimensions. Understanding relational context is vital for SSU, as the salience and interpretation of actions, expressions, and environmental cues often depend on the relationship (e.g., a gaze between colleagues versus lovers). It is also central to SSR, where mental state inferences are modulated by relational roles (e.g., a critical comment from a mentor versus a stranger implies different states). For SDP, predictions of future behaviors and counterfactual reasoning are shaped by relational norms and history (e.g., conflict resolution strategies differ between siblings and business rivals).
Figure 1: Overview of SIV-Bench, showing its diverse videos spanning various social interactions and sample QAs for three task dimensions: Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP), along with their fine-grained sub-tasks.
SIV-Bench comprises 2,792 video clips sourced from TikTok and YouTube. Its video corpus ensures robust evaluation by encompassing diverse genres (such as daily life recordings, movie clips, sports footage, and animation), varied presentation styles (including first-person, third-person, phone calls, and solo multi-role sketches) and content reflecting multiple linguistic and cultural backgrounds (like English, Japanese, and Spanish). A dedicated question-answer (QA) pipeline, combining large language models with human verification, produces 8,728 high-quality multiple-choice questions (MCQs). These MCQs are designed to target specific facets within our three core assessment dimensions. Furthermore, to systematically assess the impact of linguistic cues, SIV-Bench incorporates audio tracks and provides videos under three distinct subtitle conditions: the original version (origin), a version with transcribed and translated dialogue added (+sub), and a version with all original on-screen text removed (-sub), as illustrated in Figure 4
In the experiments, we evaluate leading commercial(e.g., Gemini-2.5-Pro [9], o4-mini [33]), open-source (e.g., InternVL3 [63], Qwen2.5-VL [3]), and specialized video models (e.g., LLaVA-OneVision [23], LLaVA-Video [59]). The results show clear performance differences across the three dimensions. Models perform relatively well in SSU but struggle with SSR, where Relation Inference proves especially challenging. Common failure patterns include confusion between primary and secondary relationships, misleading contextual cues, weak commonsense reasoning, and missed perceptual details. SDP also poses difficulties, though top models handle counterfactuals relatively well. Our study also underscores the general importance of linguistic cues, as transcribed subtitles consistently aid overall comprehension while their removal typically hinders performance.
Our key contributions are as follows: 1) We propose a novel analytical framework that structurally decomposes the complex task of multimodal social interaction understanding and reasoning into three core, interrelated dimensions, each further detailed into fine-grained sub-tasks. 2) We introduce SIV-Bench, a new video benchmark specifically curated for the analysis and comprehension of complex real-world social interactions. SIV-Bench comprises 2,792 real-world video clips representing 14 distinct social relationship types, and features 8,792 high-quality question-answer pairs generated through a human-LLM collaborative pipeline. 3) Our comprehensive experiments on diverse MLLMs reveal the limitations in their current capacity for deep human social understanding. These findings offer crucial insights to direct future research toward advancing artificial social intelligence.
2 SIV-Bench
This section details SIV-Bench, our novel video benchmark developed to evaluate MLLMs’ capabilities in understanding and reasoning within social interaction scenarios. Figure 1 offers some illustrative examples from SIV-Bench. The video corpus features interactions curated across 4 primary relational models, further detailed into 14 specific social relations, to ensure comprehensive coverage of real-world scenarios. The associated QA pairs are meticulously designed to probe model capabilities across 3 major task categories and 10 fine-grained aspects, targeting understanding, reasoning, and prediction in social contexts. Figure 2 shows the construction pipeline of SIV-Bench, including video collection (Section 2.1), QA generation and filtering (Section 2.2). Finally, to contextualize its contributions, SIV-Bench is also compared with other relevant benchmarks and datasets (Section 2.3).
Figure 2: The SIV-Bench construction pipeline, detailing the data collection process (left), and the QA generation & filtering process (right) with human-LLM collaboration. In the diagram,   blocks indicate content (like keywords, video and QA) generation steps,   blocks represent validation or modification stages, and   trapezoids signify a filtering and removal phase.
2.1 Video Collection
(a) Relation type distribution
(b) Video length distribution
(c) Language word-cloud
Figure 3: Video statistics for SIV-Bench: (a) Distribution of social relation types. (b) Distribution of video lengths, with an average of 32.49s. (c) Word cloud illustrating language diversity.
Firstly, we utilize GPT-4o-mini [20] to generate comprehensive search keywords for each of the 14 relationship types (the word-clouds are shown in Figure 7), specifically including terms associated with varying degrees of intimacy, both positive (e.g., "love", "encourage") and negative (e.g., "conflict", "fight"). Leveraging these keywords, we conduct targeted searches and download initial video candidates from TikTok and YouTube platforms using Python libraries such as TikTokApi and yt-dlp, yielding approximately 5000 raw video clips. Each video is then manually reviewed by the authors to ensure it contains clearly observable and meaningful social interactions. Videos are excluded if they do not depict clear social interaction (e.g., a vlogger speaking directly to the camera without interacting with others), if the interaction context does not fit within a set of well-defined interpersonal scenarios (e.g., an interview setting with scripted dialogue), or if the dominant interaction is difficult to identify due to the presence of multiple overlapping social dynamics (e.g., a large multi-generational family posing for a group photo). These criteria are designed to ensure that each included video primarily features one interpretable and coherent type of interpersonal interaction, allowing for more consistent analysis of social behaviors across the dataset.
The final SIV-Bench comprises 2,792 curated video clips, with statistics shown in Figure 3. The collection showcases a rich diversity in social relationships (Figure 3(a)), where Communal Sharing interactions are the most represented category, with parent-child scenarios alone accounting for 18% of the videos. The other three relational models are also well represented, ensuring broad interpersonal coverage. In terms of duration (Figure 3(b)), clips average 32.49 seconds.
Figure 4: Illustration of the three subtitle conditions applied in SIV-Bench: ‘Origin’ (original video with existing on-screen text), ‘-Subtitle’ (original text removed), and ‘+Subtitle’ (transcribed and translated dialogue added).
While most fall in the 10–20 second range, the dataset spans a wide distribution, including many short clips (under 10 seconds) and a significant number over 60 seconds. Linguistically (Figure 3(c)), English serves as the predominant language, yet SIV-Bench incorporates a multicultural dimension by including content in various other languages such as Spanish, Filipino, Korean, and Japanese, thereby reflecting a greater diversity of real-world social scenarios [30, 48]. More details about video diversity (like genre and style) are shown in Appendix A.2.
To evaluate how different forms of textual information affect MLLMs’ understanding of social interactions (Figure 4), we implement specific subtitle processing methods. Many original videos (‘Origin’) contain embedded on-screen text that often serves as scene descriptors or keywords (e.g., the ‘Buy and Sell’ text overlay). To focus on visual and auditory cues, we create a ‘-Subtitle’ version by removing such original textual overlays using video-subtitle-remover [56]. Conversely, to provide full access to spoken dialogue, we generate a ‘+Subtitle’ version. We employ Whisper-large-v3 [38] for audio transcription of the dialogue, and then use GPT to translate these transcriptions into English, ensuring consistent and high-quality subtitles (e.g., ‘He sold us his iPhone 12 with 256GB,’ as shown in ‘+Subtitle’).
2.2 QA Composing
Our Question-Answer (QA) composition process begins with generating an initial set of diverse QA pairs for each video, leveraging the capabilities of Gemini-2.0-Flash [15] which supports full video input. The prompts are carefully designed to elicit questions that span a broad range about social interactions, forming the basis for our SSU, SSR, and SDP evaluation dimensions. SSU focuses on descriptive observations of scenes, actions, and attributes; SSR targets inferences about intents, emotions, and relations; and SDP involves factual/counterfactual predictions. This stage typically yielded an average of 10 versatile QA pairs per video, designed to cover these core dimensions. Then we employ GPT-4o-mini to generate four distractors for each QA pair. These distractors are crafted to be contextually relevant to the question while remaining clearly distinguishable from the correct answer. We find that separating QA creation and distractor generation into two stages significantly improves quality and reduces ambiguity over generating all options simultaneously by a single model.
For quality assurance and to establish a gold standard, all generated QAs then undergo a rigorous consensus validation step [7, 8, 2]. We utilize three expert models (Gemini-2.0-Flash, Gemini-2.0-Pro, and GPT-4o-mini) to independently answer all questions. The QAs on which these models agree constitute our primary high-confidence set (Subset 1). The remaining QAs, where model agreement is not achieved, are flagged for further review or specialized handling (Subset 2). The following parts detail the subsequent processing strategies for each subset:
Subset 1: For Subset 1, which comprises QAs where expert models initially reached consensus (potentially indicating a need to increase challenge or verify video dependency), we implement a two-step filtering process. First, GPT-4o-mini is utilized to answer these questions without any video input. Questions that are answered correctly under this condition are removed. While this is a stringent criterion, as the model might occasionally guess correctly without sufficient information, it serves as an effective filter to ensure the questions necessitate video-based understanding, a worthwhile trade-off given our large initial pool of generated QAs. Second, the remaining questions are then scored for difficulty by GPT-4o-mini on a 1-to-5 scale. From this pool, we select the top 1-2 highest-scoring (i.e., most difficult) questions per video. Following these automated filtering stages, the authors conduct a final cross-validation review of these selected QAs to ensure their overall quality, relevance, and appropriate difficulty level. This multi-stage process ultimately yields 3,273 high-quality QAs.
Figure 5: Detailed statistics of Question-Answer pairs in SIV-Bench, showing the distribution across the 10 fine-grained sub-tasks.
Subset 2: We recruit a team of 20 human annotators to verify the non-consensus items. Each QA-pair is reviewed by at least two people. We only retain questions for which all reviewing annotators independently and unanimously select the same answer option, which is then established as the ground truth. This human validation process results in 3,096 QAs. Additionally, these annotators create novel and challenging questions along with corresponding answers. They follow specific instructions designed to target diverse abilities across the 10 fine-grained tasks. The questions are intentionally crafted to be difficult, requiring detailed understanding or complex reasoning. Appendix B.3 provides full details on our human annotation process, including demographic statistics on the 20 annotators, the guidelines for them, and the custom interface utilized. For these 2,359 newly human-generated QAs, plausible distractors are then efficiently created using GPT-4o-mini.
To ensure the quality and consistency of QA pairs, all curated QAs undergo a final automated refinement stage, where a GPT-4o-mini processes both correct answers and distractors. The goal is to standardize linguistic style and normalize option lengths. This minimizes superficial cues that models might exploit, improves test accuracy, and reduces variation from annotators or earlier generation steps. All the prompts and details in this section can be found in Appendix B.1 and B.2.
Finally, SIV-Bench contains a total of 8,728 QA pairs. Their distribution across core assessment dimensions and fine-grained tasks is illustrated in Figure 5. SRR accounts for the largest share, with 4,951 QAs, highlighting the benchmark’s focus on reasoning in social interactions. This is followed by SDP with 2,664 QAs, and SSU with 1,113 QAs. Within these categories, sub-tasks like Intent Inference (2,132 QAs) and Factual Prediction (1,449 QAs) show especially high concentrations, supporting a rich evaluation of MLLM capabilities. A detailed statistical analysis of the SIV-Bench QA pairs, confirming their structural balance and diversity (e.g., option length, answer distribution, question complexity, interrogative types), is provided in Appendix B.4 with Figure 17.
2.3 Comparison with Previous Datasets & Benchmarks
Table 1 highlights the key differences between SIV-Bench and previous datasets and benchmarks. The earlier datasets that approach social interaction primarily through the lens of social relations, such as DialogRE [58] (text-only), PIPA [45] (image-only), and ViSR [26], are often restricted to the singular task of relation recognition or limited by single modalities. They also tend to lack fine-grained task diversity and face scalability challenges due to manual annotation.
Currently, most general video understanding benchmarks (like VideoVista [25] and MVBench [24]) lack a focus on social interaction; others (like MLVU [61] and Video-Bench [32]), while touching upon it, do not fully cover social relations. Even Social-IQ 2.0 [51], which concentrates on this area, has limitations in task diversity and dynamic reasoning. In contrast, SIV-Bench is built on original data, combines manual and automatic annotations, and is one of the few benchmarks to provide subtitle and audio information, thereby supporting richer multimodal social reasoning.
Table 1: Comparison of various benchmarks. It includes several aspects: total number of items (#Items, representing the number of videos, dialogues, images, etc.), number of QA pairs (#QAs), annotation method (Anno., M/A means manually/automatic manner), and the types of tasks included (Task Types: SU for Scene Understanding, SR for State Reasoning, DP for Dynamics Prediction). Note that Task Types here refer to general task categories, which include both social and physical scenarios. The table also shows whether each benchmark includes Multi-Person Interaction, covers Various Relations, is based on newly collected data (Original Collection), and provides subtitle/audio information (S.A.)
Benchmark #Items #QAs Anno. Task Types Multi-Person Various Original S.A. SU SR DP Interaction Relations Collection Social Relation Inference DialogRE [58] 1,788 - M - PIPA [45] 37,107 - M - ViSR [26] 8,000 - M General Video Understanding and Reasoning VideoMME [13] 900 2,700 M MLVU [61] 1,730 3,102 M VideoVista [25] 894 24,906 A Social-IQ 2.0 [51] 1,000 6,000 M Perception Test [34] 11,620 38,000 M MVBench [24] 3,641 4,000 A Video-Bench [32] 5,917 17,036 A&M EgoSchema [28] 5,063 5,063 A&M SIV-Bench 2,792 8,728 A&M
3 Experiments
3.1 Settings
We evaluate a diverse set of closed- and open-source MLLMs on SIV-Bench, including Gemini-2.0/2.5-Flash [15, 14], Gemini-2.5-Pro [9], GPT-4o [20], o4-mini [33], Qwen2.5-VL-7B/72B-Instruct [3], mPLUG-Owl3 [57], InternVL3-8B/78B [63], LLaVA-OneVision [23], and LLaVA-Video [59]. Evaluations are conducted using VLMEvalKit [10], with models generating responses under their default settings. Models capable of direct video processing (e.g., Gemini) are fed raw videos, while image-only models (e.g., o4-mini) receive 16 uniformly sampled frames as input.
We use a standardized prompt (Figure 18) across all models, instructing them to return both the answer letter (e.g., ‘A.’) and the full answer text. However, we observe variations in model compliance: for instance, some models (e.g., InternVL-8B and mPLUG-Owl3) occasionally tend to output only the letter, whereas others (e.g., LLaVA-Video) may provide only the text. To robustly handle these diverse output formats and ensure accurate answer parsing, we implement a two-stage matching procedure. First, we check the model’s output for a correctly formatted option letter. If absent, we use a similarity-based comparison between the model’s raw output and the full text of all answer options to identify its selection. Model performance is then measured by the accuracy of these inferred responses, as shown in Table 2. The rest of this section provides a detailed analysis.
3.2 Main Results
Table 2: Evaluation results of MLLMs on SIV-Bench. Scores are reported for SSU, SSR, SDP, and Overall performance, detailing the impact of subtitle conditions: ‘origin,’ ‘+sub’ (added transcribed dialogue), and ‘-sub’ (subtitles removed). All scores are presented as accuracy percentages (%).
Models Params Social Scene Understanding Social State Reasoning Social Dynamics Prediction Overall origin + sub - sub origin + sub - sub origin + sub - sub origin + sub - sub Open-source MLLMs mPLUG-Owl3 7B 66.32 66.21 66.46 62.36 62.19 61.33 65.19 66.30 65.22 63.79 64.01 63.22 LLaVA-OneVision 7B 62.00 62.13 62.55 63.72 64.78 61.66 64.87 65.32 62.14 63.73 64.40 62.14 LLaVA-Video 7B 68.89 69.13 69.10 62.08 61.37 60.09 63.50 63.84 62.46 63.18 63.75 61.66 Qwen2.5-VL-7B-Instruct 7B 69.51 69.30 68.89 62.65 61.84 61.04 64.18 64.89 63.90 65.01 65.13 63.53 InternVL3-8B 8B 73.02 72.58 72.81 62.72 63.06 61.20 65.33 65.95 65.25 66.14 66.28 65.35 Qwen2.5-VL-72B-Instruct 72B 84.83 85.15 83.46 70.23 70.54 69.58 74.88 74.49 74.11 74.25 74.77 73.54 InternVL3-78B 78B 82.16 83.54 82.35 67.45 67.91 66.87 73.14 73.46 72.21 72.16 72.70 71.56 Closed-source MLLMs o4-mini - 86.77 86.90 86.33 66.87 67.25 66.10 72.86 72.45 72.11 72.30 72.43 71.59 GPT-4o - 86.94 87.34 86.29 68.28 68.58 67.73 74.82 75.80 74.05 73.76 74.29 73.12 Gemini-2.0-Flash - 86.54 86.35 86.46 67.73 68.07 66.49 73.14 73.79 72.42 72.75 73.27 71.65 Gemini-2.5-Flash - 88.56 88.84 87.32 68.12 69.09 67.25 74.67 74.97 73.05 73.67 73.82 72.53 Gemini-2.5-Pro - 90.67 90.88 90.59 71.44 71.78 70.20 75.28 75.96 74.27 76.03 76.50 75.14
Overall Performance. The evaluation of various models on SIV-Bench reveals distinct performance tiers, with closed-source models generally setting the current state-of-the-art. Notably, Gemini-2.5-Pro emerges as the top-performing model, achieving the highest overall accuracy of 76.50% when provided with subtitles. Other prominent closed-source models, such as GPT-4o and Gemini-2.5-Flash, also demonstrate strong overall performance, consistently scoring above 73%. These results indicate the advanced capabilities of leading proprietary systems.
Table 3: Average relative performance change (%) of all models when adding (‘+sub’) or removing (‘-sub’) subtitles, compared to the ‘origin’.
Condition SSU (%) SSR (%) SDP (%) Overall (%) + sub 0.178 0.234 0.482 0.447 - sub -0.300 -1.176 -0.878 -0.878
Following these, open-source Multi-modal Large Language Models (MLLMs) show considerable capabilities, although a performance gap still exists. Among them, models with a larger number of parameters exhibit significantly better results; for example, Qwen2.5-VL-72B-Instruct achieves an overall score of 74.77% with subtitles, markedly outperforming its 7B counterpart (65.13% with subtitles), even superior to the closed-source models except for Gemini-2.5-Pro. Similarly, InternVL-78B performs substantially better than InternVL-8B. While these larger open-source models demonstrate strong potential and achieve commendable overall scores, there remains room for improvement to match the leading closed-source alternatives.
Subtitle Influence. Our analysis of subtitle influence (average relative changes shown in Table 3) shows that while explicit subtitles benefit MLLMs, their impact varies across tasks. SSU, which relies on direct visual cues like actions and environments, is least affected by subtitle changes, with minimal performance shifts when subtitles are added (+sub, +0.207%) or removed (-sub, -0.333%), suggesting that visual perception alone provides sufficient grounding for this task. In contrast, the more inferential tasks—SSR and SDP—show greater sensitivity to linguistic input. Transcribed dialogue (+sub) improves SDP (+0.654%) by clarifying narrative flow and communicative intent, while removing all subtitles (-sub) most impacts SSR (-1.806%), where understanding social states becomes harder without dialogue. These findings indicate that the linguistic cues from dialogue play an important role in understanding and reasoning about multi-modal social interactions.
Task-Specific Performance. SSU, focused on recognizing visible elements, is typically the easiest, with models achieving their highest scores here. However, this isn’t always the case—for instance, LLaVA-OneVision sometimes performs better on SSR or SDP. In SSU, stronger models—especially large-scale ones, whether closed-source or high-parameter open-source—consistently outperform smaller models. This performance gap, particularly pronounced in this foundational visual task, likely reflects the superior perceptual capabilities of larger models due to greater capacity and more extensive training [1]. Following SSU, SDP tends to be the next most tractable, as it involves reasoning about interactions and predicting outcomes. SSR, which requires inferring unobservable social states like emotions and intentions, remains the most difficult. Even the top models score lowest here, underscoring the challenge of capturing implicit psychological dimensions.
Figure 6: Radar chart illustrating the comparative performance of evaluated MLLMs across the 10 fine-grained sub-tasks within SIV-Bench.
To further break down model capabilities, a radar chart (Figure 6) visualizes performance across 10 fine-grained tasks in original videos (with ‘+sub’ and ‘-sub’ detailed in Appendix C.2). Stronger models, particularly closed-source ones like the Gemini series and GPT-4o, form the outer ring, clearly outperforming smaller open-source models. In SSU, the performance gap is especially evident in Action Recognition (AR) and Facial Expression Recognition (FER), reflecting the stronger models’ ability to capture subtle visual cues. SSR presents greater challenges: while models perform moderately on Intent Inference (II) and Emotion Inference (EI), they struggle with more abstract tasks like Relation Inference (RI), which often mark the lowest points on the radar. This illustrates the difficulty of moving from surface-level cues to deeper social reasoning. In SDP, performance improves slightly, likely because tasks such as Factual Prediction (FP) and Counterfactual Prediction (CP) tap into social commonsense learned from language data. Most models—except a few like LLaVA-OneVision—perform better on CP than FP, possibly due to the more explicit reasoning cues provided by hypothetical framing. Some failure cases are illustrated in Figure 19, 20, 21.
3.3 Failure Pattern Analysis of Relation Inference
Accurate Relation Inference (RI) is crucial for understanding social interactions. The difficulty lies in synthesizing diverse cues—scene context, human appearance, verbal and non-verbal expressions, physical proximity, and interaction dynamics—all of which contribute to subtle and implicit relational patterns. Our fine-grained analysis (Figure 6) highlights RI as a consistent performance bottleneck. Given its significance and the persistent challenges across models, we conduct a deeper qualitative analysis to examine common failure patterns:
Lack of Primary-Secondary Relation Differentiation. In multi-relational scenarios, MLLMs sometimes identify a valid but secondary relationship instead of the most salient one. For instance, in a video where employees celebrate their boss’s birthday, the model misclassifies the situation as <colleagues>, overlooking the underlying hierarchical relationship that defines the interaction. Similarly, when a coach angrily reprimands players for their poor performance, the model incorrectly assigns the relationship as <teammates>, failing to capture the coach-athlete authority structure that governs the scene. These errors suggest a deficiency in relational prioritization and in the ability to infer which social bond is most significant within a complex interpersonal context.
Scene and Human-Induced Misleading Cues. MLLMs are easily misled by both scene settings and human behaviors. They may over-rely on environmental cues, assigning stereotypical roles based on location without validating actual interactions. For instance, in a classroom setting where two travel agency employees co-present, it incorrectly classifies them as <teacher-student> instead of <colleagues>. Similarly, human language can mislead the model: a couple jokingly referring to each other as "best brother" is misclassified as <siblings>, despite strong visual evidence—such as different ethnicities and intimate gestures—indicating a romantic relationship.
Deficiency in Commonsense Reasoning. In several instances, the model’s decisions are inconsistent with basic commonsense knowledge. Two <friends> wearing different sports team uniforms but showing friendly behavior are misclassified as <teammates>, disregarding the fact that such uniforms typically signify competition rather than collaboration. Similarly, a store owner photographing a bustling crowd is mistaken for a colleague of the customers, whereas commonsense suggests a <transactional> relationship. These mistakes highlight the model’s limited ability to integrate external knowledge about social conventions and situational functions, emphasizing a shortfall in commonsense reasoning and real-world social understanding.
Ignoring Key Perceptual Details. Although MLLMs can capture many surface features, they will overlook subtle but critical details necessary for correct role recognition. Age differences and facial features crucial to distinguishing between <siblings> and <parent-child> relationships are sometimes ignored. Additionally, important textual clues embedded in subtitles are sometimes overlooked, such as the line "POV: you work with your coworker," which results in <colleague> not being correctly assigned. These errors suggest insufficient fine-grained perception, selective attention to critical multimodal information, and a failure to integrate minor yet decisive cues into final role judgments.
4 Related Works
Video Benchmarks for MLLMs. The evaluation of rapidly advancing Multimodal Large Language Models (MLLMs) in the video domain relies on diverse benchmarks that collectively incorporate a rich set of features for comprehensive model assessment. Many, such as MLVU [61], video-MME [13], MVbench [24], MMBench-Video [11], and Video-Bench [32], offer multiple tasks—ranging from question answering and summarization to complex reasoning—across diverse video fields including movies, egocentric recordings, and general web content. This variety aims to assess a broad spectrum of capabilities, from foundational visual understanding and detail analysis (as targeted by Perception Test [34]) (explored by VideoVista [25]) to higher-level comprehension and decision-making. A significant trend within these evaluations is the robust focus on long-form video understanding, with benchmarks like CINEPILE [40], Egoschema [28], and LVbench [50] specifically challenging models on temporal reasoning and extended context retention, while others like TempCompass [27] provide specialized assessments of critical nuanced aspects such as detailed temporal dynamics. While these benchmarks thoroughly assess diverse video comprehension capabilities, SIV-Bench offers a more specialized focus, concentrating on the intricate understanding and reasoning required to interpret dynamic social interactions among multiple people.
Evaluating Social Intelligence in AI Systems. The pursuit of artificial social intelligence is inherently multifaceted, with research progressing from early dataset construction to contemporary evaluations of large language models (LLMs). Initial efforts typically offered valuable yet modality-constrained or component-specific resources, targeting areas such as action recognition (e.g., Kinetics [22], UCF101 [43]), facial expression and contextual emotion inference (e.g., AffectNet [31], MELD [35]), and social relation recognition from static images or text (e.g., PIPA [45], DialogRE [58]). As the field has evolved, benchmarks like Social-IQ 2.0 [51] have expanded the focus toward interpersonal reasoning and social commonsense. With the rise of LLMs, more specialized evaluations have emerged. ToMATO [41] probes Theory of Mind through role-play-based interactions. SOTOPIA [62] and its extension SOTOPIA-$\pi$ [49] introduce interactive environments for developing agentive social skills. EgoSocialArena [18] offers a first-person perspective benchmark targeting both cognitive and behavioral intelligence. Social Genome [29] provides multimodal, grounded traces of social reasoning from video data, while MM-SOC [21] evaluates MLLMs on multimodal social media content. Despite these advances, most existing benchmarks focus on isolated competencies. A comprehensive evaluation framework that holistically assesses MLLMs’ understanding of complex social interactions—spanning perception, social state reasoning, and dynamics prediction in real-world video contexts—remains lacking. SIV-Bench is introduced to fill this gap.
5 Conclusion
In this work, we introduce SIV-Bench, a novel video benchmark designed to rigorously evaluate Multimodal Large Language Models’ capabilities in understanding and reasoning about complex social interactions. Featuring 2,792 diverse videos categorized by foundational social relationship models and 8,792 meticulously generated question-answer pairs, SIV-Bench structures evaluation around three core dimensions: Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP), further detailed into ten fine-grained sub-tasks. Our comprehensive experiments reveal that while leading MLLMs adeptly handle SSU, they still struggle with SDP and SSR, where Relation Inference emerges as an acute bottleneck. Moreover, our study underscores the critical impact of transcribed dialogue on comprehending complex social interactions.
Although SIV-Bench provides a strong evaluation framework, its current scale suggests room for future expansion into a larger training corpus with broader social contexts and more diverse content. The multiple-choice format also offers a basis for future extensions into interactive or generative evaluation. By identifying current model strengths and weaknesses, SIV-Bench offers timely insights and a foundation for advancing socially intelligent AI.
Appendix A Video Details
A.1 Video Collection
To compile a rich and diverse video corpus for SIV-Bench that reflects a wide array of social interactions, we first identify 14 key social relationship types (e.g., parent-child, friends, colleagues, service interactions). For each of these types, we employ GPT-4o mini to generate a comprehensive set of search keywords. This strategy is designed to ensure broad coverage in our video collection from platforms like YouTube and TikTok, capturing various interaction categories (such as cooperation, conflict, and support), forms (including verbal discussions, non-verbal expressions, and shared activities), and differing levels of intimacy or formality inherent to each relationship. Figure 7 visualizes these tailored keyword sets as word clouds for each of the 14 relationship types, illustrating the breadth of concepts targeted. This targeted yet diverse keyword generation is crucial for assembling a video dataset that truly represents the complexity of human social life.
(a) Grandparent-Child
(b) Parent-Child
(c) Siblings
(d) Couple
(e) Friends
(f) Boss-Employee
(g) Coach-Player
(h) Teacher-Student
(i) Caregiver-Recipient
(j) Classmates
(k) Teammates
(l) Colleague
(m) Transactional
(n) Service
Figure 7: Word clouds of GPT-generated keywords used for sourcing videos across 14 distinct social relationship types (e.g., (a) Grandparent-Child, (b) Parent-Child, (e) Friends, (f) Boss-Employee). Keywords were designed to capture diverse interaction scenarios, forms, and intimacy levels from platforms like YouTube and TikTok.
A.2 Video Diversity
The SIV-Bench video corpus is intentionally diverse in both subject matter and presentation to ensure comprehensive model evaluation. This genre diversity, illustrated in Figure 8, includes a wide array of video types such as candid Daily Life recordings, scripted Movie Clips, dynamic Sports Replays, illustrative Commercials, and non-photorealistic Animated Videos. Complementing this, Figure 9 showcases varied visual presentation styles, ranging from traditional Third-Person perspectives and immersive First-Person (POV) footage to interactions depicted via Phone Call interfaces and creative Solo Multi-Role Sketches. This multifaceted diversity in video content and form provides a robust and varied testbed.
Figure 8: Examples illustrating the diverse video genres in SIV-Bench, including (from left to right) Daily Life recordings, Movie Clips, Sports Replays, Commercials, and Animated Videos, ensuring a broad range of social interaction contexts.
Figure 9: Examples of diverse video presentation styles featured in SIV-Bench, including (from left to right) Third-Person views, First-Person perspectives, Phone Call interfaces, and Solo Multi-Role Sketches.
Appendix B QA Details
B.1 Automatically QA Generation
The prompt shown in Figure 10 guides the Gemini in generating diverse and challenging Question-Answer (QA) pairs that form the foundation of SIV-Bench. Its design ensures these QAs effectively cover our core evaluation dimensions.
Figure 10: The prompt used to guide Gemini for the initial generation of question-answer pairs.
To create the multiple-choice options for each QA pair, we utilize a dedicated prompt for distractor generation, as detailed in Figure 11. This prompt instructs GPT to produce four unique distractor options based on the given question and its corresponding correct answer. Crucially, the prompt emphasizes that these distractors must be reasonable and clearly distinguishable from the correct answer, while also maintaining consistency with the correct answer in terms of sentence length, language style, and grammar. Furthermore, the prompt specifies a JSON output format for the list of distractors to ensure seamless programmatic integration. This structured approach to distractor generation is essential for constructing high-quality, fair multiple-choice questions that minimize biases arising from option formatting and robustly test model comprehension.
Figure 11: The prompt used to generate distractors for each QA pair.
B.2 Automatically QA Filtering
To ensure that the QA pairs in SIV-Bench genuinely require understanding of the video content, particularly for Subset 1, we employ a video-agnostic filtering step. Figure 12 presents the prompt used to instruct GPT for this purpose. The central directive of this prompt is for the model to select the most likely answer from the provided options using only the textual content of the question and the options themselves, without any access to the video. The prompt also enforces a strict output format, requiring only the letter of the chosen answer, and instructs the model to make a random choice if it deems the information insufficient (though the primary test is correct identification based on text alone). If a QA pair is answered correctly under this video-absent condition, it indicates that the question may be solvable through world knowledge, commonsense reasoning, or cues within the text of the QA itself, rather than necessitating specific information from the video. Such QAs are subsequently removed from the candidate pool to enhance the benchmark’s focus on video-grounded social understanding.
Figure 12: Prompt for the video-agnostic filtering of Question-Answer pairs.
To further refine the SIV-Bench dataset, particularly in selecting challenging items for Subset 1, we employ GPT-4o-mini to assign a difficulty score to each Question-Answer (QA) pair. Figure 13 presents the prompt designed for this task. It directs the model to evaluate QA difficulty on a 1 (Very Easy) to 5 (Very Hard) scale, assuming the QA pertains to a social interaction video. The prompt outlines key assessment criteria, including the subtlety of relevant cues, the complexity of reasoning involved, and the depth of social understanding needed to arrive at the correct answer. This automated difficulty scoring enables us to systematically identify and prioritize QAs that demand more sophisticated comprehension, thereby enhancing the overall rigor of our benchmark.
Figure 13: Prompt used to score the difficulty of SIV-Bench QA pairs on a 1-to-5 scale, based on defined criteria such as cue subtlety, reasoning complexity, and depth of social understanding required.
To classify each QA pair in SIV-Bench, we use a GPT model guided by a structured prompt (see Figure 14). The prompt defines our three main assessment dimensions—SSU, SSR, and SDP—and their sub-tasks. The model analyzes each QA and outputs its corresponding sub-task, ensuring consistent and precise categorization across the benchmark.
Figure 14: Prompt used to instruct LLM for the final classification of Question-Answer pairs into one of the 10 fine-grained sub-tasks under SSU, SSR and SDP.
B.3 Human Annotation
Annotator Information. For the human annotation tasks integral to SIV-Bench, we recruit a team of 20 annotators. This cohort consists of 12 female and 8 male individuals, with an average age of 27 years. All annotators are well-educated and demonstrate proficiency in English, which is the primary language for instructions and annotations. To ensure fair compensation for their detailed work, annotators are remunerated at an hourly rate of \$4 USD.
Annotator Guidelines. To maintain consistency and high quality in human annotation, particularly for the creation of new challenging Question-Answers (QAs) for Subset 2, annotators are provided with a comprehensive set of guidelines, as illustrated in Figure 15. These instructions detail the process for responding to existing multiple-choice questions and, crucially, guide annotators in formulating one new challenging question per video. For this QA creation task, annotators are directed to ensure their questions primarily test one of our three core assessment dimensions: Social Scene Understanding (SSU), Social State Reasoning (SSR), or Social Dynamics Prediction (SDP). The guidelines, including example questions, also emphasize that new QAs must be demanding, clearly worded, and require a deep understanding of the video content rather than superficial observation.
Annotation Process. All human annotation tasks for SIV-Bench, including the review of existing Question-Answers (QAs) and the generation of new QAs, are conducted using a custom-designed web interface. Figure 16 provides a representative example of this interface, which allows annotators to view the social interaction video, respond to provided multiple-choice questions, and submit their own newly authored questions and answers based on the video content and provided guidelines. This platform ensures a standardized environment for all human annotation contributions.
Figure 15: Screenshot of the guidelines provided to human annotators, detailing the tasks of answering multiple-choice questions and creating new challenging questions about Social Scene Understanding (SSU), Social State Reasoning (SSR), and Social Dynamics Prediction (SDP).
Figure 16: Example of the web-based interface used by human annotators for watching videos, answering provided multiple-choice questions, and authoring new Question-Answer pairs for SIV-Bench.
B.4 QA statistics
To ensure the quality and balance of our SIV-Bench Question-Answer (QA) pairs, we conducted a statistical analysis of their structural properties, as summarized in Figure 17. This analysis examines several aspects: the average word count per multiple-choice option (Figure 17(a)) is relatively consistent across options A through E, minimizing length-based cues. The distribution of correct answers (Figure 17(b)) is fairly uniform across the five options, preventing positional bias. Furthermore, the overall question length (Figure 17(c)) peaks at around 11 words but shows a broad range, indicating variability in question complexity. Finally, an analysis of the first word in questions (Figure 17(d)) reveals a diverse set of interrogative types, led by ’what,’ ’why,’ and ’which,’ reflecting a variety of reasoning challenges posed to the models.
(a) Word count per option
(b) Correct answer distribution
(c) Question word count
(d) First word in questions
Figure 17: Statistical analysis of SIV-Bench Question-Answer (QA) pairs. (a) average word count consistency across answer options, (b) distribution of correct answers among options, (c) overall question length distribution by word count, and (d) frequency of common first words in questions.
Appendix C Experiments Details
C.1 Settings Details
This section provides further details on our experimental setup for evaluating MLLMs on SIV-Bench. Figure 18 displays the standardized prompt templates employed for model evaluations, with distinct versions tailored to models based on their input capabilities. For models that process sequences of images, the "PROMPT for frames input" (Figure 18, Top) informs the MLLM that it will receive a set of uniformly sampled frames from a video in chronological order. For models capable of direct video processing, the "PROMPT for videos input" (Figure 18, Bottom) is used. Both prompts clearly instruct the MLLM on its role, the task of answering multiple-choice questions based on the provided visual input, the expected JSON-like format for organizing answers (providing the exact text of the chosen option for each question), and a strict directive to avoid any extraneous text such as explanations or conversational remarks. This standardized, yet input-adaptive, prompting approach ensures consistency in task presentation across different model architectures.
For all evaluations, the specific inference parameters used for each model—such as temperature, top-p, or maximum new tokens—are adopted from their default configurations as provided within the VLMEvalKit [10] framework. This adherence to default settings aims to reflect the out-of-the-box capabilities of these models and ensure fair comparability. The experiments are conducted on two primary compute clusters. Cluster 1, utilized for evaluating the largest open-source models (Qwen2.5-VL-72B-Instruct and InternVL-78B), is equipped with an AMD EPYC 7642 48-Core Processor and 4x NVIDIA A100 GPUs. The total runtime for the reported experiments on this cluster is approximately 3 days. Cluster 2, used for the remaining models, consists of an Intel(R) Xeon(R) Platinum 8369B CPU @ 2.90GHz and 8x NVIDIA RTX 3090 GPUs. The cumulative runtime for experiments on this cluster is approximately 2 days. It should be noted that the overall research project, including preliminary testing on earlier dataset versions and exploratory experiments not included in the final results, involved a greater amount of compute time than the specific durations reported for the final benchmark evaluations.
Figure 18: Standardized prompt templates used for evaluating MLLMs on SIV-Bench. Separate prompts are shown for models that process (Top) uniformly sampled frames and (Bottom) direct video input.
C.2 Results Details
Table 4: Comparative accuracy (%) of MLLMs on the 10 fine-grained sub-tasks in SIV-Bench (Origin), grouped by SSU, SSR, and SDP dimensions.
Models SSU SSR SDP Overall AR FER HAI EP II EI AI RI FP CP mPLUG-Owl3 62.36 68.35 59.50 70.61 62.81 65.08 61.17 60.13 56.98 74.98 63.79 LLaVA-OneVision 59.55 65.60 71.90 59.34 70.36 55.79 56.74 61.18 70.06 58.68 63.73 LLaVA-Video 60.11 63.76 71.07 78.41 64.17 61.22 56.54 60.05 65.19 61.48 63.18 Qwen2.5-VL-7B-Instruct 58.71 64.22 70.25 81.26 68.15 68.21 59.96 56.58 64.99 63.21 65.01 InternVL3-8B 64.05 72.48 67.77 82.47 70.08 70.32 61.77 53.03 60.90 70.62 66.14 Qwen2.5-VL-72B-Instruct 86.52 85.78 80.99 84.01 71.72 75.81 67.81 66.99 74.76 77.04 74.25 InternVL3-78B 82.30 81.65 82.65 82.16 73.78 73.15 63.18 61.02 72.15 74.32 72.16 o4-mini 86.80 90.83 77.69 87.26 72.42 67.46 65.80 64.73 72.05 73.83 72.30 GPT-4o 88.48 84.40 80.99 88.67 70.45 78.02 66.20 64.65 74.41 75.31 73.16 Gemini-2.0-Flash 90.73 83.03 85.12 85.21 74.11 67.03 67.20 64.41 69.87 77.04 72.15 Gemini-2.5-Flash 91.29 85.32 87.60 88.20 74.48 67.04 67.61 64.97 71.86 78.03 73.67 Gemini-2.5-Pro 93.54 91.74 89.26 88.08 75.14 76.94 69.62 67.80 72.01 79.18 76.03
Table 5: Comparative accuracy (%) of MLLMs on the 10 fine-grained sub-tasks in SIV-Bench (+sub), grouped by SSU, SSR, and SDP dimensions.
Models SSU SSR SDP Overall AR FER HAI EP II EI AI RI FP CP mPLUG-Owl3 62.60 68.55 59.69 70.83 63.03 65.30 61.36 60.33 57.17 75.22 64.01 LLaVA-OneVision 60.22 66.25 72.59 60.01 71.03 56.47 57.41 61.88 70.71 59.37 64.40 LLaVA-Video 60.65 64.30 71.61 78.98 64.75 61.75 57.12 60.63 65.76 62.07 63.75 Qwen2.5-VL-7B-Instruct 58.81 64.34 70.35 81.34 68.29 68.34 60.08 56.68 65.10 63.35 65.13 InternVL3-8B 64.21 72.60 67.90 82.65 70.27 70.44 61.89 53.17 61.01 70.75 66.28 Qwen2.5-VL-72B-Instruct 87.02 86.31 81.52 84.53 72.23 76.34 68.32 67.51 75.28 77.55 74.77 InternVL3-78B 82.87 82.21 83.20 82.69 74.31 73.66 63.73 61.57 72.69 74.90 72.70 o4-mini 86.91 90.94 77.81 87.39 72.54 67.62 65.90 64.88 72.18 74.00 72.43 GPT-4o 89.60 85.51 82.12 89.83 71.53 79.17 67.38 65.78 75.53 76.45 74.29 Gemini-2.0-Flash 91.83 84.12 86.22 86.34 75.22 68.15 68.31 65.50 70.98 78.14 73.27 Gemini-2.5-Flash 91.42 85.44 87.72 88.34 74.63 67.19 67.73 65.09 72.00 78.19 73.82 Gemini-2.5-Pro 94.02 92.25 89.77 88.52 75.61 77.43 70.09 68.25 72.43 79.66 76.50
Table 6: Comparative accuracy (%) of MLLMs on the 10 fine-grained sub-tasks in SIV-Bench (-sub), grouped by SSU, SSR, and SDP dimensions.
Models SSU SSR SDP Overall AR FER HAI EP II EI AI RI FP CP mPLUG-Owl3 61.77 67.76 58.92 70.03 62.26 64.52 60.60 59.58 56.41 74.42 63.22 LLaVA-OneVision 57.98 64.00 70.30 57.73 68.77 54.20 55.15 59.58 68.45 57.10 62.14 LLaVA-Video 58.59 62.24 69.51 76.89 62.64 59.68 55.00 58.52 63.71 59.93 61.66 Qwen2.5-VL-7B-Instruct 57.23 62.72 68.78 79.79 66.62 66.71 58.48 55.10 63.53 61.76 63.53 InternVL3-8B 63.26 71.69 66.99 81.66 69.30 69.57 60.96 52.22 60.12 69.86 65.35 Qwen2.5-VL-72B-Instruct 85.83 85.08 80.28 83.31 71.02 75.11 67.09 66.32 74.03 76.29 73.54 InternVL3-78B 81.70 81.06 82.09 81.56 73.22 72.56 62.57 60.39 71.55 73.71 71.56 o4-mini 86.10 90.11 76.97 86.57 71.73 66.75 65.09 63.99 71.35 73.12 71.59 GPT-4o 88.43 84.37 80.94 88.64 70.39 77.96 66.20 64.60 74.38 75.22 73.12 Gemini-2.0-Flash 90.24 82.50 84.63 84.73 73.55 66.55 66.73 63.89 69.35 76.52 71.65 Gemini-2.5-Flash 90.17 84.20 86.45 87.07 73.34 65.87 66.46 63.85 70.71 76.90 72.53 Gemini-2.5-Pro 92.69 90.84 88.36 87.15 74.24 76.08 68.72 66.90 71.10 78.28 75.14
C.3 Failure Cases
This section presents several illustrative failure cases. We focus on examples from Gemini-2.0-Flash, a strong closed-source model, to highlight that even advanced MLLMs can fail on nuanced aspects of social perception, prediction, and reasoning. These examples are categorized by our primary assessment dimensions: SSU, SSR, and SDP, and are intended to offer concrete instances for future research and model development.
Figure 19 illustrates instances where Gemini-2.0-Flash fails on SSU tasks, which require accurate perception of explicit visual elements. (a) In Action Recognition, the model incorrectly identifies the man’s gesture as "He crosses his arms tightly" instead of the correct "He raises one eyebrow slightly", missing a subtle but distinct facial action. (b) For Environment Perception, when asked about the weather, the model failed to capture the details of the characters in the scene wearing thick scarves and down jackets to infer that the correct answer was "cold", but instead wrongly chose "wet". (c) In Facial Expression Recognition, the model describes the expression as "A mischievous smile" rather than the correct "A stoic glare", misinterpreting the nuanced facial expression display. (d) For Human Attribute Identification, concerning the child’s clothing, the model selects "A dress" instead of the correct "A set of pajamas", failing to correctly identify common apparel.
Figure 20 presents failure cases of Gemini-2.0-Flash on SSR tasks, which involve inferring unobservable mental states and relationships. (e) In Intent Inference, when a woman says "do you understand?" to a boy who bullies her son in an angry tone, it is to teach him a lesson and warn him not to bully her son again, not for "discourage any defiance", because in fact no child much younger than her can form defiance against her. (f) For Emotion Inference, this employee is happy instead of scared after leaving because he successfully deceives the boss into giving him a vacation. (g) In Attitude Inference, the coworker is dissatisfied and disappointed with the cashier’s nervousness, panic and even physical reactions when seeing female customers. This could also be seen from his subsequent warning to the cashier not to do so anymore. (h) For Relation Inference, we present case studies on the failure patterns of the four common models listed in the main text in this task.
Figure 21 highlights errors made by Gemini-2.0-Flash in SDP tasks, which require predicting future events or reasoning about hypothetical scenarios. (i) In Factual Prediction, when asked if the person in the black shirt would be satisfied with the workers’ work, since the two of them have already reached an agreement with smiles at the end of the video, it could be inferred that the answer was "yes", but the model chooses another answer. (j) For Counterfactual Prediction, the video shows the dance interaction between a mother and her son. The question raised is what would happen if one of them had more dance experience. This can be inferred from the positive and relaxed interaction between the two in the video. The most likely answer is "The more experienced dancer leads and adapts movements". For example, a son leads his mother to learn dancing happily, rather than "The less experienced dancer hesitates and struggles to keep up". They have a good relationship, and the probability of negative performance like "hesitates" and "struggles" is lower.
Figure 19: Examples of failure cases in Social Scene Understanding (SSU) tasks, including errors in Action Recognition, Environment Perception, Facial Expression Recognition, and Human Attribute Identification.
Figure 20: Examples of failure cases in Social State Reasoning (SSR) tasks, highlighting difficulties in Intent Inference, Emotion Inference, Attitude Inference, and Relation Inference.
Figure 21: Examples of failure cases in Social Dynamics Prediction (SDP) tasks, covering both Factual Prediction and Counterfactual Prediction.
Appendix D Broader Impact
SIV-Bench is designed to foster positive advancements in artificial social intelligence, potentially leading to more empathetic, context-aware, and collaborative AI systems for beneficial applications such as assistive technologies, improved human-AI teaming, and richer content understanding. However, enhancing AI’s grasp of social dynamics also presents risks. These capabilities could be misused for sophisticated manipulation, disinformation, or invasive surveillance, and unaddressed biases in data could be amplified, leading to inequitable outcomes. We offer SIV-Bench as a research tool to transparently assess MLLM capabilities and limitations in the social domain, thereby encouraging the community to proactively consider these ethical challenges and develop robust safeguards alongside continued innovation in social AI.